How much pure acid should e added to 51 cc of 12% acid to increase the concentration of 15%?
Answer by mananth(16946)

percent ---------------- quantity
Acid solution 12 ----------------51 cc
Pure acid 100 ----------------x cc
Mixture 15 ----------------51 + x cc

12*51+100x=15(51+x)
612+100 x=765+15x
100x-15x=765-612
85 x = 153
/ 85
x= 1.8 cc of Pure acid